Solve the equations by the use of the operator D







D²y -6Dy +9y= e^3x + e^-3x




1
Expert's answer
2021-07-01T17:31:41-0400
"(D-3)^2(y)=e^{3x} + e^{-3x}"

"(D-3)^2=e^{3x}D^2e^{-3x}"

Then


"e^{3x}D^2e^{-3x}(y)=e^{3x} + e^{-3x}"

"D^2e^{-3x}(y)=e^{3x} + e^{-3x}"

We have the general solution od homogeneous differential equation


"y_h=(A+Bx)e^{3x}"


"L(D)\bigg[\dfrac{e^{kx}}{L(k)}\bigg]=e^{kx}"

"L(D)=D^2-6D+9"

Since


"L(3)=(3)^2-6(3)+9=0"

we use


"y_1=\dfrac{1}{1}(\dfrac{1}{2!}x^2)e^{3x}=\dfrac{1}{2}x^2e^3x"


"y_2=\dfrac{e^{-3x}}{L(-3)}"

"L(-3)=(-3)^2-6(-3)+9=36"

The particular solution is


"y_p=\dfrac{1}{2}x^2e^{3x}+\dfrac{1}{36}e^{-3x}"

The genral solution of nonhomogeneous differential equation is


"y=(A+Bx)e^{3x}+\dfrac{1}{2}x^2e^{3x}+\dfrac{1}{36}e^{-3x}"